Assume that you take a flask, evacuate it to remove all the air, and finds its mass to be 478.1 g. You then fill the flask with argon to a pressure of 2.15 atm and reweigh it. What would the balance read (in grams) if the flask has a volume of 7.35 L and the temperature is 20.0 oC?


1
Expert's answer
2021-04-25T07:37:09-0400

Pv=nRT

2.15×7.35=n×8.32×20

n= 0.0950

Therefore mass=0.095×40=3.8g

But original mass=478+3.8=481.9gm